OpenStudy (sloppycanada):

Check my work? If a wheel with a radius of 30 inches spins at a rate of 500 revolutions per minute, find the approximate linear velocity in miles per hour. 89 MPH?

OpenStudy (sloppycanada):

@thushananth01 If you're not too busy

OpenStudy (anonymous):

V = r.w w = 2.pi/T w = 2.pi.f v = r .2.pi.f = 30000pi inches per min 89.24 mph Yes you are right!

OpenStudy (anonymous):

\[\frac{\ \pi*2\ *30\ *500\ *60}{12*5280}=\frac{625 \pi }{22}=89.2498 \text{ mph} \]